# Lifelong Graph Summarization with Neural Networks: 2012, 2022, and a Time Warp

### v1

@@ 4. Experimental Apparatus
In this section, we introduce our dataset and pre-processing, followed by a description of the experimental procedure. Afterward, we present how we optimize our hyperparameters. Finally, we present performance and forgetting measures for the evaluation of lifelong graph learning.
4.1. DyLDO Dataset in 2012 and 2022
@@
The DyLDO dataset [^2] http://km.aifb.kit.edu/projects/dyldo/data/ was introduced by Käfer et al. (2012). It consists of weekly snapshots of the web. Starting from a seed list of about $90.000$ Uniform Resource Identifiers (URIs), it samples data from the web to create a snapshot. A snapshot consists line-based, plain text format for encoding Resource Description Framework (RDF) graphs (Gavin Carothers, 2014). The first snapshot was created on 6 May, 2012. The data contains some faulty and duplicate lines, which we removed following Blasi et al. (2022).
In our experiments, we use the following snapshots: The snapshots from 6 May, 2012 to 8 July, 2012 and the snapshots from 25 September, 2022 to 27 November, 2022. We have 20 snapshots in total, ten from each year. We use the ten snapshots from 2012 since that is when the sampling process started. The different numbers of EQCs in Figures 2(a) and 2(d) show that the dataset has much changed over the ten years. Thus, we use other snapshots from 2022, as well.
For each snapshot, we create the labels of the vertices by calculating their EQCs according to the summary model. We remove each vertex in a snapshot with a degree higher than $100$ for $M_{AC2}$ to be able to train on enough samples in a reasonable time for $M_{AC2}$. We also exclude them from the validation and test set. This has only a small effect on our dataset. For example, the number of EQCs removed for the summary $M_{AC2}$ of the first snapshot of DyLDO in 2012 is less than $5\%$, i. e., down from $288,418$ to $275,993$ EQCs.
4.2. Dataset Analysis
@@
Figure 2 shows that the EQCs in each year change from snapshot to snapshot. We observe that the EQCs of the $2$-hop summary change more than the EQCs of the $1$-hop. Both changes underlie the same snapshot change, meaning that the $2$-hop summary changes more than the $1$-hop summary in the same sequence of snapshots. In general, there are more different EQCs in the $2$-hop summaries computed from the snapshots of 2022 than in 2012. The numbers of EQCs in the $1$-hop summaries are similar between the two years.
@@
(b) $1$-hop, May–July 2012
(c) $2$-hop, May–July 2012
(d) September–November 2022
(e) $1$-hop, Sept.–Nov. 2022
@@
Figure 2. Detailed analyses of the DyLDO snapshots from 2012 and 2022 used in our experiments. Figures (a) and (d) show the absolute numbers of EQCs regarding the snapshots for a $1$-hop versus $2$-hop model in 2012 and 2022, respectively. The EQCs resemble the number of classes the neural networks are trained on. Figures (b), (c), (e), and (f) show the changes of the EQCs (addition, deletion, and recurring) in comparison to the previous snapshot.
4.3. Procedure
GNNs and baselines
We use a standard MLP with one hidden layer and ReLU-activation functions with dropout as a baseline. For $M_{AC1}$, we use an MLP because all information is contained in the feature vector. We have two cases for $M_{AC2}$, one with encoded edge information of the neighbors of a vertex and one without. In the first case, we only use GraphSAINT because the other neural networks do not use neighborhood information during inference. In the second case, we use all neural networks, i. e., GraphSAINT, MLP, and Graph-MLP. We train all networks with $100$ batches, each containing up to $1,000$ vertices (as described in Section 3.5), based on pre-experiments.
Experiments Summarizing DyLDO Snapshots from 2012 and 2022
We perform our experiments twice, once for our snapshots of $2012$ and once for our snapshots of $2022$. We incrementally train our network on each snapshot in chronological order. The vertices of each snapshot are divided into $2\%$ validation data and $5\%$ testing data. The remaining $93\%$ of the data is used for sampling the training data. Because each snapshot is very large, we can only train on parts of it.
For the first snapshot of each year, the neural networks are initialized with an input layer and an output layer with the size of the number of predicates and the number of classes, respectively. Then, the networks are trained on the snapshot for $100$ iterations with maximum batch size $1,000$. Each iteration has its own sampled batch. After training, we increase the input and output layer, if the next snapshot has more classes or predicates. The extended neural network is trained on the next snapshot for $100$ iterations. We repeat this process until we reach the last snapshot.
After we trained our networks, we tested each of the networks on all snapshots. We do this to measure the backward and forward transfer of the different networks (see Section 4.5).
@@ 4.3. Procedure
In this experiment, we take the neural network trained on the 2012 snapshots and test it on the first snapshot in 2022. We contrast this reuse of a “10-year-old” network with a variant where the network is trained on the first snapshot of 2022. Finally, we compare these two “time warped” networks with a newly initialized network that is only trained on the first snapshot in 2022. This experiment shows whether reusing old networks of 2012 is helpful despite the time gap of ten years, compared to starting training from scratch in 2022.
4.4. Hyperparameter Optimization
@@
For the hyperparameter search, we apply grid search on the parameter ranges from Blasi et al. (2022). We tune the learning rate on the values $\{0.1,0.01,0.001\}$, dropout in $\{0.0,0.2,0.5\}$, and hidden layer size $\{32,64\}$ for the GNNs. For Graph-MLP, we optimize the weighting coefficient $\alpha\in\{1.0,10.0,100.0\}$, the temperature $\tau\in\{0.5,1.0,2.0\}$, learning rate in $\{0.1,0.01\}$, and the hidden layer size $\{64,256\}$. For MLP, we use a hidden layer size of $1,024$ following prior works (Galke and Scherp, 2021; Blasi et al., 2022). We use Adam as an optimizer.
@@
The best configurations depend on the summary model, namely $M_{AC1}$ and $M_{AC2}$. For the summary model $M_{AC1}$, we use an MLP with a hidden layer of size $1,024$, a dropout of $0.5$, and a learning rate of $0.01$. For $M_{AC2}$ with edge information, we use GraphSAINT with two hidden layers of size $32$, learning rate $0.1$, and no dropout. For $M_{AC2}$ with no edge information, we use an MLP with one hidden layer of size $1,024$, dropout $0.5$, and learning rate $0.01$. Furthermore, we use Graph-MLP with layer size $64$, dropout $0.2$, learning rate $0.01$, $\alpha=1$, and $\tau=2$. Finally, for GraphSAINT we use one hidden layer with hidden dimension $64$, learning rate $0.1$, and no dropout. Details can be found in Appendix A.
4.5. Measures
@@
We assess vertex classification performance using accuracy. For lifelong learning, we base our analysis on the measures from Lopez-Paz and Ranzato (2017) and Chaudhry et al. (2018).
@@
We train a network for each snapshot and test it on all other snapshots. Thus, for each of the years 2012 and 2022, we compute a result matrix $R\in\mathbb{R}^{T\times T}$ with $T=10$. Each entry $R_{i,j}$ is the test accuracy of the network on task $\mathcal{T}_{j}$ after observing the last sample from task $\mathcal{T}_{i}$. That means the network is trained on task $\mathcal{T}_{1}$ to $\mathcal{T}_{i}$ and tested on $\mathcal{T}_{j}$. We use these matrices to calculate the following measures for each year, to measure the backward and forward transfer of the networks. The average accuracy (ACC) a network achieves on all tasks $i$ after being trained on the last task $T$ is given by
$ACC=\frac{1}{T}\sum^{T}_{i=1}R_{T,i}\,.$
The backward transfer (BWT), which indicates how much knowledge of previous tasks the network keeps after being trained on the last task $T$ is given by
$BWT=\frac{1}{T-1}\sum^{T-1}_{i=1}R_{T,i}-R_{i,i}\,.$
The forward transfer (FWT), which shows how much knowledge of a task is reused for the next one, is defined as [^3] Different from Lopez-Paz and Ranzato (2017) who use the test accuracy $\bar{b}_{i}$ for the task $i$ at random initialization as a comparison, we use the performance $R_{i,i}$.
$FWT=\frac{1}{T-1}\sum_{i=2}^{T}R_{i-1,i}-R_{i,i}\,.$
We use the matrices to calculate further measures. The diagonal $R_{i,i}$ resembles the test accuracy for task $i$ after it is learned. Furthermore, $R_{i,1}$ is the test accuracy of the test data of the first task after $i$ tasks have been learned, i. e., the network is trained on the first $i$ snapshots and then tested in the first snapshot. Let $\alpha_{\mathrm{ideal}}=\max_{i\in\{1,\dots,T\}}R_{i,i}$ be the maximal test accuracy of the network trained up to task $i$ and then evaluated on task $i$. We define $\alpha_{i,\mathrm{all}}=\frac{1}{T}\sum^{T}_{t=1}R_{i,t}$ as the average accuracy of a network trained up to task $i$. With this, we can define the following measures (Kemker et al., 2018):
$\Omega_{\mathrm{base}}$ $=\frac{1}{T-1}\sum^{T}_{i=2}\frac{R_{i,1}}{\alpha_{\mathrm{ideal}}}$
$\Omega_{\mathrm{new}}$ $=\frac{1}{T-1}\sum^{T}_{i=2}R_{i,i}$
$\Omega_{\mathrm{all}}$ $=\frac{1}{T-1}\sum^{T}_{i=2}\frac{\alpha_{i,\mathrm{all}}}{\alpha_{\mathrm{ideal}}}\,.$
@@
The first snapshot is the base test set, and the best accuracy $\alpha_{\mathrm{ideal}}$ in the matrix is the optimum. With $\Omega_{\mathrm{base}}$, we measure how a network performs on the base test set after training on other sets compared to the optimum. $\Omega_{\mathrm{new}}$ shows how a network performs on a snapshot after being trained on it. $\Omega_{\mathrm{all}}$ measures how a network performs on all previous snapshots compared to the optimum.
Finally, we use the accuracy matrix for computing the forgetting measure (Chaudhry et al., 2018):
$F_{k}=\frac{1}{k-1}\sum^{k-1}_{j=1}f_{j}^{k}\,,$
@@
where $f_{j}^{k}$ is the forgetting on task $j$ after the network is trained up to task $k$ and is computed as:
$f_{j}^{k}=\max_{\ell\in\{1,\dots,k-1\}}R_{\ell,j}-R_{k,j}\,.$
5. Results
@@
We describe the results for the vertex classification and analyze the lifelong learning measures. Afterward, we look into the forgetting of the networks. Finally, we investigate the ten-year time warp.
@@
(c) Graph-MLP ($2$-hop)
(d) GraphSAINT ($2$-hop)
(e) GraphSAINT ($2$-hop and edges)
Figure 3. Accuracies for snapshots trained from May to July 2012.
@@
(c) Graph-MLP ($2$-hop)
(d) GraphSAINT ($2$-hop)
(e) GraphSAINT ($2$-hop and edges)
Figure 4. Accuracies for snapshots trained from September to December 2022.
The classification accuracies for the snapshots in 2012 and 2022 are shown in Figures 3 and 4, respectively. In general, the networks perform well when they are tested on task $i$ after they were trained up to snapshot $i$.
Regarding 2012, the performance of the MLP for $M_{AC1}$ is much higher than for $M_{AC2}$. The difference between these two MLPs is between $0.25$ and $0.41$. Furthermore, the best performance of a network is observed when tested on the first snapshot independent of the training snapshots. For 2022, the results are similar to the experiments in 2012 but with overall lower accuracies. The MLP accuracies for $M_{AC1}$ are again much higher than those for $M_{AC2}$, but the difference is between $0.33$ and $0.53$. The best result for each snapshot is achieved by the network that has been trained on it. In contrast to 2012, the results on the first snapshot are not better than in the subsequent snapshots.
Table 1. Lifelong learning measures applied to the different network and summaries of each year.
@@
2022 GraphSAINT $2$-hop 0.298 -0.276 -0.311 0.426 0.543 0.421
2022 GraphSAINT $2$-hop and edges 0.269 -0.273 -0.282 0.431 0.514 0.426
The lifelong learning measures applied to our measured accuracies are shown in Table 1. It shows that in 2012 and 2022, no network independent of the used summary model or information achieved a positive forward or backward transfer in our setting. For the other measures, the MLP applied on $M_{AC1}$ achieves the best results, and the results for $M_{AC2}$ are similar, independent of the network and the year.
Network Forgetting
Figure 5. The progress of the forgetting in 2012.
Figure 6. The progress of the forgetting in 2022.
@@
The forgetting of the networks in 2022, as shown in Figure 6, differs from the ones in 2012. The highest forgetting for the $2$-hop networks is right in the beginning, and the lowest forgetting value is achieved by the MLP ($1$-hop) after being trained on the 16 October snapshot. In general, the forgetting values are higher in 2022 than in 2012. Nevertheless, after the networks are trained on the last snapshot, they have similar forgetting values again.
@@
Table 2. Accuracy after training on the task from September 25, 2022. Trained from scratch (no time warp, no TW) or initialized with weights from July 8, 2012 (/w TW).
Neural network Summary model /w TW no TW
@@
GraphSAINT $2$-hop 0.565 0.572
GraphSAINT $2$-hop and edges 0.526 0.515
We show the results of reusing a neural network from 2012 to classify vertices in 2022 in Table 2. We call this setting a ten-year time warp (TW). Retraining the last network from 2012 on the first snapshot in 2022 gives the same results as training a network from scratch in 2022. Table 2 shows that the MLP with $M_{AC1}$ performs a bit better in the time warp. For $M_{AC2}$, the time warp makes the performance of the neural network better in two of four cases, and worse in the other cases. Reusing the last network from 2012 without retraining on the first 2022 snapshot resulted in very low accuracy – below $0.03$ in all cases (not shown in the table).
6. Discussion

### v2

@@ IV Experimental Apparatus
In this section, we introduce and analyze our dataset, and describe our experimental procedure. Afterward, we present how we optimize our hyperparameters. Finally, we present performance and forgetting measures for the evaluation of lifelong graph learning.
IV-A Datasets: Weekly DyLDO Snapshots from 2012 and 2022
@@
We use the weekly snapshots of the DyLDO web crawl as foundation [7]. Starting from a seed of $90,000$ Uniform Resource Identifiers, it samples data from the web to create a snapshot. A snapshot consists line-based, plain text encoding of an Resource Description Framework (RDF) graph [29].
We create two datasets, using the following sequences of snapshots: The ten snapshots from 6 May, 2012 to 8 July, 2012 (the first ten crawls) and the ten snapshots from 25 September, 2022 to 27 November, 2022. Figures 2(a) and 2(b) show the number of EQCs in each snapshot.
For each snapshot, we create the labels of the vertices by calculating their EQCs according to the summary model. For $M_{\mathrm{AC2}}$ we remove all vertices with a degree over $100$ from the training, validation, and test sets to speed up the experiments.
This has only a small effect on the number of EQCs: for example, in the first snapshot of DyLDO in 2012 only $12,425$ out of $288,418$ EQCs are removed, which is less than $5\%$.
IV-B Dataset Analysis
@@
Figure 2 shows how the EQCs in each year change from snapshot to snapshot. We observe that the EQCs of the $2$-hop summary change more than those of the $1$-hop summary. Both changes underlie the same snapshot modification, indicating that the $2$-hop summary changes more than the $1$-hop summary in the same sequence of snapshots. In general, there are more EQCs in the $2$-hop summaries in 2022 than in 2012. The numbers of EQCs in the $1$-hop summaries are similar between the two years.
@@
(b) September–November 2022
(c) $1$-hop, May–July 2012
(d) $1$-hop, Sept.–Nov. 2022
(e) $2$-hop, May–July 2012
@@
Figure 2: Detailed analyses of the DyLDO snapshots used in our experiments. Figures (a) and (b) show the number of unique EQCs per snapshot for a $1$-hop versus $2$-hop model. Figures (c)–(f) show the changes of the EQCs (addition, deletion, and recurring) compared to the previous snapshot.
IV-C Procedure
Training of GNNs and Baselines
Based on pre-experiments, we train all networks with $100$ batches, each containing up to $1,000$ vertices. Our baseline is a standard MLP with one hidden layer and ReLU-activation functions with dropout. For $M_{\mathrm{AC1}}$, we use an MLP because all information is contained in the feature vector. We have two cases for $M_{\mathrm{AC2}}$, one with encoded edge information and one without. In the first case, we only use GCN because the other neural networks do not use neighborhood information during inference. In the second case, we use all neural networks, i. e., GCN, MLP, and Graph-MLP.
Lifelong Graph Learning for Summarizing the DyLDO Snapshots
We perform the lifelong graph learning experiments once for the $2012$ and once for the $2022$ snapshots. We train the networks on the snapshots in chronological order. We split the vertices of each snapshot into $2\%$ validation, $5\%$ testing and sample the training data from the remaining $93\%$.
For the first snapshot of each year, the neural networks are initialized with an input size of number of predicates and output size of the number of classes. Then, the networks are trained on the snapshot for $100$ iterations with maximum batch size $1,000$. After training, we increase the input and output size if the next snapshot has more classes (EQCs) or predicates. The extended neural network is trained on the next snapshot for $100$ iterations. We repeat this process until the last snapshot. After training the networks on a snapshot, we evaluate them on all snapshots of the same year. We do this to calculate the lifelong learning measures (see Section IV-E).
@@ IV-C Procedure
In this experiment, we compare three networks on the first snapshot of 2022. One “10-year-old” network is only trained on the 2012 snapshots, one is initialized from the old network but also trained on the 2022 snapshot, and one which is trained from scratch just on the 2022 snapshot. This experiment shows whether reusing old networks of 2012 is helpful despite the time gap of ten years, compared to starting training from scratch in 2022.
IV-D Hyperparameter Optimization
@@
For the hyperparameter search, we apply grid search. We tune the learning rate on the values $\{0.1,0.01,0.001\}$, dropout in $\{0.0,0.2,0.5\}$, and hidden layer size $\{32,64\}$ for the GNN. For Graph-MLP, we optimize the weighting coefficient $\alpha\in\{1.0,10.0,100.0\}$, the temperature $\tau\in\{0.5,1.0,2.0\}$, learning rate in $\{0.1,0.01\}$, and the hidden layer size $\{64,256\}$. For MLP, we use a hidden layer size of $1,024$ following prior works [30, 6]. We use Adam as an optimizer.
@@
The best configuration depends on the summary model. Details can be found in our appendix. For the summary model $M_{\mathrm{AC1}}$, we use an MLP with a hidden size of $1,024$, a dropout of $0.5$, and a learning rate of $0.01$. For $M_{\mathrm{AC2}}$ with edge information, we use GCN with two hidden layers of size $32$, learning rate $0.1$, and no dropout. For $M_{\mathrm{AC2}}$ with no edge information, we use an MLP with one hidden layer of size $1,024$, dropout $0.5$, and learning rate $0.01$. We use Graph-MLP with hidden size $64$, dropout $0.2$, learning rate $0.01$, $\alpha=1$, and $\tau=2$. GCN uses one hidden layer of dimension $64$, learning rate $0.1$, and no dropout.
IV-E Measures
@@
We assess vertex classification performance using test accuracy, a well-known measure in graph neural networks like GCN [13]. For lifelong learning, we use the measures from Lopez-Paz and Ranzato [10] and Chaudhry et al. [11].
@@
We train a network for each snapshot and test it on all other snapshots. Thus, for each of the years 2012 and 2022, we compute a result matrix $R\in\mathbb{R}^{T\times T}$ with $T=10$. Each entry $R_{i,j}$ is the test accuracy of the network on task $\mathcal{T}_{j}$ after observing the last sample from task $\mathcal{T}_{i}$. That means the network is trained on tasks $\mathcal{T}_{1}$ to $\mathcal{T}_{i}$ and tested on $\mathcal{T}_{j}$. We use these matrices to calculate the following measures for each year, to measure the backward and forward transfer of the networks. The average accuracy (ACC) a network achieves on all tasks $\mathcal{T}_{i}$ after being trained on the last task $\mathcal{T}_{T}$ is given by $ACC=\frac{1}{T}\sum^{T}_{i=1}R_{T,i}$. The backward transfer (BWT), which indicates how much knowledge of previous tasks the network keeps after being trained on the last task $T$ is given by $BWT=\frac{1}{T-1}\sum^{T-1}_{i=1}(R_{T,i}-R_{i,i})$. The forward transfer (FWT), which shows how much knowledge of a task is reused for the next one, is based on Lopez-Paz and Ranzato [10] as $FWT=\frac{1}{T-1}\sum_{i=2}^{T}(R_{i-1,i}-R_{i,i})$.
@@
We define the ideal performance as $\alpha_{\mathrm{ideal}}=\max_{i\in\{1,\dots,T\}}R_{i,i}$. We define $\alpha_{i,\mathrm{all}}=\frac{1}{T}\sum^{T}_{t=1}R_{i,t}$ as the average accuracy of a network trained up to task $\mathcal{T}_{i}$. With this, we can define the following measures, as in [31]: $\Omega_{\mathrm{base}}=\frac{1}{T-1}\sum^{T}_{i=2}\frac{R_{i,1}}{\alpha_{\mathrm{ideal}}}\text{, }\Omega_{\mathrm{new}}=\frac{1}{T-1}\sum^{T}_{i=2}R_{i,i}\text{, and }\Omega_{\mathrm{all}}=\frac{1}{T-1}\sum^{T}_{i=2}\frac{\alpha_{i,\mathrm{all}}}{\alpha_{\mathrm{ideal}}}\,.$ The first snapshot is the base test set, and the best accuracy $\alpha_{\mathrm{ideal}}$ in the matrix is the optimum. $\Omega_{\mathrm{base}}$ measures how a network performs on the base test set after training on other sets compared to the optimum. $\Omega_{\mathrm{new}}$ shows how a network performs on a snapshot after being trained on it. $\Omega_{\mathrm{all}}$ measures how a network performs on all previous snapshots compared to the optimum.
@@
Finally, we compute the forgetting measure [11]: $F_{k}=\frac{1}{k-1}\sum^{k-1}_{j=1}f_{j}^{k}$, where $f_{j}^{k}$ is the forgetting on task $j$ after the network is trained up to task $k$ and is computed as: $f_{j}^{k}=\max_{\ell\in\{1,\dots,k-1\}}(R_{\ell,j}-R_{k,j})$.
V Results
@@
We describe the results for the vertex classification and the lifelong learning measures. Finally, we show the results of the ten-year time warp.
@@
Figure 3: Accuracies for snapshots trained from May to July 2012. The accuracies for Graph-MLP ($2$-hop), GCN ($2$-hop), and GCN ($2$-hop and edges) are omitted because they are similar to the values of the MLP ($2$-hop).
@@
Figure 4: Accuracies for snapshots trained from September to December 2022. Graph-MLP ($2$-hop), GCN ($2$-hop), and GCN ($2$-hop and edges) are omitted because they are similar to the values of the MLP ($2$-hop).
The classification accuracies for the snapshots in 2012 and 2022 are shown in Figures 3 and 4. In general, the networks perform best on task $i$ after they were trained up to snapshot $i$.
Regarding 2012, the performance of the MLP for $M_{\mathrm{AC1}}$ is much higher than for $M_{\mathrm{AC2}}$. The difference between these two MLPs is between $0.25$ and $0.41$. Furthermore, the best performance of a network is observed when tested on the first snapshot independent of the training snapshots. For 2022, the results are similar to the experiments in 2012 but with overall lower accuracies. The MLP accuracies for $M_{\mathrm{AC1}}$ are again much higher than those for $M_{\mathrm{AC2}}$. The difference is between $0.33$ and $0.53$. The best result for each snapshot $i$ is achieved by the network that has been trained on snapshots $1$ to $i$. In contrast to 2012, the results on the first snapshot are not better than in the subsequent snapshots.
TABLE I: Lifelong learning measures applied to the different networks and summaries of each year.
@@
2022 GCN $2$-hop $0.298$ $-0.276$ $-0.311$ $0.426$ $0.543$ $0.421$
2022 GCN $2$-hop and edges $0.269$ $-0.273$ $-0.282$ $0.431$ $0.514$ $0.426$
Lifelong Learning Measures
Figure 5: The progress of the forgetting in 2012.
Figure 6: The progress of the forgetting in 2022.
The lifelong learning measures are shown in Table I. It shows that in 2012 and 2022, no network with any summary model or information achieved a positive forward or backward transfer in our setting. For the other measures, the MLP applied on $M_{\mathrm{AC1}}$ achieves the best results, and the results for $M_{\mathrm{AC2}}$ are similar, independent of the network and the year.
@@
The forgetting of the networks in 2022, as shown in Figure 6, differs from the ones in 2012. The highest forgetting for the $2$-hop networks is right in the beginning, and the lowest forgetting value is achieved by the MLP ($1$-hop) after being trained on the 16 October snapshot. In general, the forgetting values are higher in 2022 than in 2012. Nevertheless, after the networks have been trained on the last snapshot, they have similar forgetting values.
@@
TABLE II: Accuracy after training on the task from September 25, 2022. Trained from scratch (no time warp, no TW) or initialized with weights from July 8, 2012 (TW).
Neural network Summary model TW no TW
@@
GCN $2$-hop $0.565$ $0.572$
GCN $2$-hop and edges $0.526$ $0.515$
We show the results of reusing a neural network from 2012 to classify vertices in 2022 in Table II. We call this setting a ten-year time warp (TW). Retraining the last network from 2012 on the first snapshot in 2022 gives the same results as training a network from scratch in 2022. Table II shows that the MLP for $M_{\mathrm{AC1}}$ performs a bit better in the time warp. For $M_{\mathrm{AC2}}$, the time warp makes the performance of the neural network better in two of four cases, and worse in the other two. Reusing the last network from 2012 without retraining on the first 2022 snapshot gives very low accuracy – always below $0.03$ (not shown in the table).
VI Discussion